OpenStudy (jjuden):

do the tables below show a quadratic variation? x: 1,2,3,4, f(x): 4,16,36,64 x: 1,2,3,4 f(x): 3,16,45,96

OpenStudy (anonymous):

are they the corresponding values ?

OpenStudy (jjuden):

idk

OpenStudy (anonymous):

if yes ..den try doing this : f(x)= ax^2 + bx+ c..... say , for x=1 , f(x)=4...den u ll get 1 eqn : 4=a+b+c....put it like dis for all 4 cases..if u get d same values for a,b,c den it is a quad function ......i think so..

OpenStudy (jjuden):

ok

OpenStudy (anonymous):

WELCOME !